OpenStudy (anonymous):

It takes 1.5 litres of sugar soap to wash a square metre of ceiling and 0.5 litres of sugar soap to wash a square metre of wall. Bianca’s room is 4 metres high and each wall is 5 metres wide. There is one window, which is 4m2 and a door, which is 9m2. How much sugar soap will Bianca need, assuming she does not wash the door or window (a room has four walls and a roof)?

OpenStudy (anonymous):

too easy!

OpenStudy (anonymous):

first calculate total area to be washed

OpenStudy (anonymous):

each wall is 5m wide * 4m high = 20m^2 now there are four walls ( excluding the ceiling for now ) 1 wall has a window on it and 1 wall has a door on it and the other 2 walls have nothing on them so: the wall with the window on it, you have to minus the area of the window from the area of the wall ! and the same thing with the door! 1 WALL WITH WINDOW AREA = 20 - 4 = 16m^2 1 WALL WITH DOOR AREA = 20 - 9 = 11m^2 2 NORMAL WALLS = 20 x 2 = 40m^2 Total area = the sum of all the 4 walls so 16 + 11 + 40 = 67m^2 Now to find the amount of sugar soap required to wash these walls with total area of 67m^2 Take 67 / 0.5 = 134 litres for the walls Now for the ceiling with a total area of 5 x 5 = 25m^2 Take 25 / 1.5 = 16.67 litres for the ceiling